A Bibliography of Publications of Luc Brun

Luc Brun L.E.R.I

Computer science Rue des Crayères

Reims France

Tel: (33).3.26.05.30.88 FAX: (33).3.26.05.30.48

E-mail: brun@leri.univ-reims.fr (Internet)

13 October 2017 Version 1.05

Abstract

This bibliography records publications of Luc Brun.

Title word cross-reference

n [FB09].

-D [FB09].

5th [BV05].

6th [MMU96].

7th [TEB09].

’96 [MMU96].

algorithm [BD97]. Amélioration [BB97b, BB94]. Analysis [BBV96]. April [BV05].

base [Bru96]. based [BV05, TEB09]. boundaries [BD97].

Color [BB97a]. Combinatorial [BK01, BK02, BK00, BMM05, BK05, BK06, FB09, BK03]. Comparison [BB97a]. computer [MMU96]. conceptual [DHB09]. construction [PB07]. Contains [BK06]. couleur [BB94, Bru96]. couleurs [BB95].

D [FB09]. defined [BpD96]. detection [DHB09]. DGCI [MMU96]. d’images [Bru96]. discrete [BD96, BpD96, MMU96]. d’un [BB97b]. dynamique [BB94, BB95].

Euclidean [BBV96]. expert [BB97b].

Fields [BK02, BK03]. first [FB09]. Framework [BK02, BK03, BMM05, PB07]. France [BV05, MMU96].

GbRPR [BV05, TEB09]. geometry [MMU96]. graph [DHB09, BV05, TEB09]. Graph-based [BV05, TEB09].

Hierarchical [BMM05]. Hierarchy [PB07].

IAPR [BV05, TEB09]. IAPR-TC-15 [TEB09]. Image [BBV96, BB96, BB97a, DHB09, BpD96]. imagery [MMU96]. images [BD96, DHB09]. Incremental [BpD96]. Inter [BBV96, BB96, BD97]. Inter-Pixel [BBV96, BB96, BD97]. international [BV05, MMU96, TEB09]. interpretation [DHB09]. Introduction [BK01]. Irregular [BK00]. Italy [TEB09].

labeling [DHB09]. Lyon [MMU96].

Maps [BB96, BD96, BpD96, BD97, BK00]. May [TEB09]. merge [BD97]. méthodes [BB94]. Methods [BB97a]. modifications [BpD96].

November [MMU96].

objects [DHB09]. Optimization [BB97a]. over-segmented [DHB09].

parition [BB95]. partition [BB94]. Paths [BBV96]. pattern [BV05, TEB09]. performances [BB97b]. Pixel [BBV96, BB96, BD97]. Poitiers [BV05]. Proceedings [BV05, TEB09, MMU96]. Pyramid [BK02, BMM05, BK03]. Pyramids [BK01, BK00, BK05, BK06, FB09].

Quantification [BB95, BB94]. Quantization [BB97a].

Receptive [BK02, BK03]. recognition [BV05, TEB09]. region [BD96]. relationships [BK06]. Representation [BB96, BD96]. representations [BV05, TEB09].

scale [PB07]. schemes [PB07].

Segmentation [BB96, Bru96, BB97b]. segmented [BD96, BpD96, DHB09]. set [PB07]. spaces [FB09]. split [BD97]. step [FB09]. système [BB97b].

TC [TEB09]. Topological [BB96, BD97]. topologique [Bru96].

unexpected [DHB09]. utilisation [BB97b].

Venice [TEB09].

watersheds [BMM05]. within [BK02, BK03, BMM05, BK05, BK06, PB07]. workshop [BV05, MMU96, TEB09].

References

Braquelaire:1994:AMQ

[BB94] J. P. Braquelaire and L. Brun. Une amélioration des méthodes de quantification de couleur par partition dynamique. In Deuxiemes journées de l’association française d’informatique graphique, pages 247–261. A.F.I.G, ????, ????, 1994.

Braquelaire:1995:QCP

[BB95] J. P. Braquelaire and L. Brun. Quantification de couleurs par parition dynamique. In Actes de la journée COULEUR ET INFOR-MATIQUE GRAPHIQUE. ????, ????, mars 1995.

Braquelaire:1996:IST

[BB96] J. P. Braquelaire and L. Brun. Image segmentation with topological maps and inter-pixel representation. To appear in Journal of Visual Communication and Image representation., May 1996.

Braquelaire:1997:COM

[BB97a] J. P. Braquelaire and L. Brun.

Comparison and optimization of methods of color image quantization. IEEE Transactions on Image Processing, 6(7):1048–1052, July 1997. CODEN IIPRE4. ISSN 1057-7149 (print), 1941-0042 (electronic).

Brun:1997:APS

[BB97b] L. Brun and J. M. Bazin.

Amélioration des performances d’un système de segmentation par l’utilisation d’un système expert. Submitted to IPMU., November 1997.

Braquelaire:1996:IPE

[BBV96] J. P. Braquelaire, L. Brun, and Anne Vialard. Inter-pixel euclidean paths for image analysis. In Miguet et al. [MMU96], pages 193–204. ISBN 3-540-62005-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 1996.

Braquelaire:1996:RRS

[BD96] J. P. Braquelaire and J.-P.

Domenger. Representation of region segmented images with discrete maps. Technical report, RR-112797 LaBRI, available on http://www.labri.u-bordeaux.fr/LaBRI/Publications/index.html, June 1996.

Brun:1997:NSM

[BD97] L. Brun and Jean Philipe Domenger. A new split and merge algorithm with topological maps and interpixel boundaries. In The Fifth International Conference in Central Europe on Computer Graphics and Visualization. ????, ????, February 1997.

Brun:2000:IPC

[BK00] Luc Brun and Walter Kropatsch. Irregular pyramids with combinatorial maps. In Francesc J. Ferri et al., editors, Advances in pattern recognition. Joint IAPR international workshops SSPR 2000 and SPR 2000, Alicante, Spain, August 30–September 1, 2000. Proceedings, volume 1876 of Lecture Notes in Computer Science, pages 256–265. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/1876/18760256.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/1876/
18760256.pdf.

Brun:2001:ICP

[BK01] Luc Brun and Walter Kropatsch. Introduction to combinatorial pyramids. Lecture Notes in Computer Science, 2243:108–128, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2243/22430108.htm;   http:
//link.springer-ny.com/link/

service/series/0558/papers/2243/
22430108.pdf.

Brun:2002:RFW

[BK02] Luc Brun and Walter G. Kropatsch.

Receptive fields within the combinatorial pyramid framework. Lecture Notes in Computer Science, 2301:92–101, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.
com/link/service/series/0558/
bibs/2301/23010092.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/2301/
23010092.pdf.

Brun:2003:RFW

[BK03] Luc Brun and Walter Kropatsch.

Receptive fields within the Combinatorial Pyramid framework. Graphical models, 65(1–3):23–42, May 2003. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic).

Brun:2005:IOW

[BK05] Luc Brun and Walter Kropatsch.

Inside and outside within combinatorial pyramids. In Brun and Vento [BV05], pages 122–131. EUR 52.00, SFR 92.00, £40.00, US$72.00.

Brun:2006:CIR

[BK06] Luc Brun and Walter Kropatsch.

Contains and inside relationships within combinatorial pyramids. Pattern Recognition, 39(4):515– 526, 2006. CODEN PTNRA8. ISSN 0031-3203 (print), 1873-5142 (electronic).

Brun:2005:HWW

[BMM05] Luc Brun, Myriam Mokhtari, and Fernand Meyer. Hierarchical watersheds within the combinatorial pyramid framework. In Eric Andres et al., editors, Discrete geometry for computer imagery. 12th international conference, DGCI 2005, Poitiers, France, April 13– 15, 2005. Proceedings, volume 3429 of Lecture Notes in Computer Science, pages 34–44. Springer-Verlag Inc., New York, NY, USA, 2005.

Brun:1996:IMS

[BpD96] L. Brun and Jean philippe Domenger. Incremental modifications on segmented image defined by discrete maps. Technical report, RR-112696 LaBRI, May 1996. Submitted.

Brun:1996:SIC

[Bru96] L. Brun. Segmentation d’images couleur à base topologique. PhD thesis, Université Bordeaux I, 351 cours de la Libération, 33405 Talence, France, December 1996.

Brun:2005:GBR

[BV05] Luc Brun and Mario Vento, editors. Graph-based representations in pattern recognition. 5th IAPR international workshop, GbRPR 2005, Poitiers, France, April 11– 13, 2005. Proceedings, volume 3434 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. EUR 52.00, SFR 92.00, £40.00, US$72.00.

Deruyver:2009:IIC

[DHB09] Aline Deruyver, Yann Hodé, and Luc Brun. Image interpretation with a conceptual graph: labeling over-segmented images and detection of unexpected objects. Artificial Intelligence, 173(14):1245– 1265, 2009. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Fourey:2009:FST

[FB09] Sébastien Fourey and Luc Brun.

A first step toward combinatorial pyramids in n-D spaces. In Torsello et al. [TEB09], pages 304–313. EUR 56.66.

Miguet:1996:DGC

[MMU96] Serge Miguet, Annick Montanvert, and Stephane Ubeda, editors. Discrete geometry for computer imagery: 6th international workshop, DGCI ’96, Lyon, France, November 13–15, 1996: proceedings, volume 1176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. ISBN 3-540-62005-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN T385 .D525 1996.

Pruvot:2007:HCS

[PB07] Jean-Hugues Pruvot and Luc Brun.

Hierarchy construction schemes within the scale set framework. In Francisco Escolano et al., editors, Graph-based representations in pattern recognition. 6th IAPR-TC-15 international workshop, GbRPR 2007, Alicante, Spain, June 11– 13, 2007. Proceedings, volume 4538 of Lecture Notes in Computer Science, pages 126–137. Springer-Verlag Inc., New York, NY, USA, 2007.

Torsello:2009:GBR

[TEB09] Andrea Torsello, Francisco Escolano, and Luc Brun, editors. Graph-based representations in pattern recognition. 7th IAPR-TC-15 international workshop, GbRPR 2009, Venice, Italy, May 26–28, 2009. Proceedings, volume 5534 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. EUR 56.66.